OpenStudy (anonymous):

In a 30°- 60°- 90° right triangle, the longer leg is cm. How long are the other two sides of the triangle?

OpenStudy (anonymous):

The longer leg is 6 sqrt 3 (:

OpenStudy (anonymous):

ratios of a 30-60-90 triangle are \(1:\sqrt{3}:2\) for short side : long side : hypotenuse

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so, 1:6 sqrt 3: 2?

OpenStudy (anonymous):

so if long slide is \(6\) then short sides is \(\frac{6}{\sqrt{3}}\)

OpenStudy (anonymous):

Gotcha. So how do I find the Hypotenuse?

OpenStudy (anonymous):

better known as \(2\sqrt{3}\)

OpenStudy (anonymous):

double the short side

OpenStudy (anonymous):

lengths are therefore \(2\sqrt{3}, 6,4\sqrt{3}\)

OpenStudy (anonymous):

btw i hope it is clear that \[\frac{6}{\sqrt{3}}=\frac{6}{\sqrt{3}}\times \frac{\sqrt{3}}{\sqrt{3}}=\frac{6\sqrt{3}}{3}=2\sqrt{3}\]

OpenStudy (anonymous):

Oh okay.. But my answers to choose from are different. The only one that comes close is, shorter leg: 6 sqrt 3, and hypotenuse: 6

OpenStudy (anonymous):

oooh because i read your problem incorrectly it says longer leg is \(6\sqrt{3}\) then it is much easier

OpenStudy (anonymous):

Hehe. (: So, is my answer right?

OpenStudy (anonymous):

to get the short side, divide by \(\sqrt{3}\) and get 6

OpenStudy (anonymous):

And then double that and get 12? Right? (:

OpenStudy (anonymous):

then the hypotenuse is twice that, namely 12 yes you got it

OpenStudy (anonymous):

Awesome! Thanks soo much :D You teach better then my teacher @satellite73

OpenStudy (anonymous):

thank you for the compliment (blush) sorry i did not read your question right the first time

OpenStudy (anonymous):

It's fine, really. Could you help me with another?